# The influence of Parker spiral on the reflection-driven turbulence

### v1

@@ 4.2 Energy Evolution and Heating
Figure 5 shows the normalized wave action energies, $\tilde{E}^{\pm}/\tilde{E}^{+}_{0}$, as functions of expansion factor $a$. Here $\tilde{E}^{+}_{0}=\tilde{E}^{+}(a=1)$ is the initial outward wave-action energy, so its decrease provides a measure of turbulent heating per (16). Plotted are HR-$\Phi_{0}\!=\!0^{\circ},4^{\circ}$; A05-$\Phi_{0}\!=\!0^{\circ},2^{\circ},5^{\circ}$, and MR-$\Phi_{0}\!=\!0^{\circ},2^{\circ},5^{\circ}$ runs spanning initial $\chi_{\rm exp,0}$ in [15, 45][$\approx$ 15,\;45]. For runs with smaller $\chi_{\rm exp,0}$ we find an initial approximate power-law decay $\tilde{E}^{+}\propto a^{-0.6}$ or $a^{-0.5}$; for larger $\chi_{\rm exp,0}$ the decay is faster with $\sim a^{-0.8}$. This discrepancy with the $a^{-1}$ limit predicted in RDT phenomenology of section˜2.4 (Eq. 15) is likely because the $a^{-1}$ decay is recovered only for sufficiently large $\chi_{\rm exp,0}$. A similar conclusion was found in Meyrand2025, who found that very large $\chi_{\rm exp}$ was needed to recover the expected $a^{-1}$ decay.
Focusing on the HR runs, the radial case displays a slow change near $a\!\sim\!15$ where $\tilde{E}^{+}$ stops decaying. We will show below that this is consistent with $\chi_{\rm exp}$ approaching unity, as discussed in the section 2. Concurrently $\tilde{E}^{-}$ grows with expansion, following a similar trend, and approaching $\tilde{E}^{+}$ as the system moves towards a quasi-balanced state. During this transition from a strongly imbalanced turbulent state to a quasi balanced sate, the nonlinear coupling weakens because $\bm{z}^{+}$ and $\bm{z}^{-}$ becomes proportional (see Fig. 3 and 4) and the turbulent heating rate drops. Eventually, as shown by Meyrand2025 we expect $\tilde{E}^{+}$ to rise as the system starts developing a very large negative residual energy,
@@
The simulations also allow a cursory examination of the influence of fluctuation amplitude on the energy evolution. Figure 5 compares (a) a high-amplitude case with $\rm A=1.0$ and (b) a lower amplitude case with $\rm A=0.5$ and similar $\chi_{\rm exp,0}$. In the lower amplitude case, the fluctuations remain nearly linearly polarized (reduced-MHD like), and follow a slower energy decay $\tilde{E}^{+}\propto a^{-0.5}$, while higher-amplitude runs, which must have larger $\delta B_{\parallel}$ to maintain spherical polarization exhibit steeper decay. One possible mechanism to explain this difference is that, the anisotropic reflection term (see Eq. 9) that seeds $\tilde{\bm{z}}^{-}$ from $\tilde{\bm{z}}^{+}$ modifies the character of the reflection once fluctuations become large. For small-amplitude, nearly transverse Alfvénic fluctuations, the reflection operator $\tilde{\mathbb{T}}=\mathrm{diag}(-1,1,1)$ leads to $\tilde{z}^{-}_{y,z}\propto-\tilde{z}^{+}_{y,z}$ with $\tilde{z}^{\pm}_{x}$ negligible. At large amplitude, however, finite $\delta B_{\parallel}$ (nonzero $\tilde{z}_{x}^{\pm}$) and the oblique mean field mix components: the 1-1 in $\tilde{\mathbb{T}}$ produces additional contributions to $\tilde{z}_{x}^{-}\propto\tilde{z}_{x}^{+}$. In other words, high amplitude fluctuations scramble the simple $\tilde{\bm{z}}^{-}\propto-\tilde{\bm{z}}^{+}$ relation, reduce the magnitude of $\langle\tilde{\bm{z}}^{+}\!\cdot\!\tilde{\bm{z}}^{-}\rangle$, and thereby allow stronger nonlinearity to persist in the PS geometry compared with the radial case. In-situ near-Sun measurements often show large normalized fluctuation amplitudes (Bale2019), the high-amplitude runs are likely more representative of observed solar wind.
@@ 4.3 Evolution of length-scales and $\chi_{exp}$
Figure 6: Evolution of correlation length $\ell$ as a function of $a$ for magnetic-field ($\bm{B_{\perp}}$) fluctuations for MR-$\Phi_{0}\!=\!0^{\circ}$ and MR-$\Phi_{0}\!=\!2^{\circ}$ runs. Each curve represents projections along the field ($\parallel$, blue), perpendicular($\perp$, brown), $\hat{e}_{T}$ direction ($\perp,T$, green), and $\hat{e}_{N}$ direction ($\perp,N$, orange). The dashed black line indicates the $a^{1}$ power-law scaling for reference. In the radial case (left panel), $\ell_{\perp}$ grows nearly linearly with $a$, consistent with expansion-driven eddy widening at early times then transitions to faster growth at late times (see Meyrand2025). In contrast, the Parker spiral geometry modifies the growth: $\ell_{\perp,T}$ saturates and remains nearly constant, while $\ell_{\perp,N}$ increases more slowly than linear.
Figure 6 shows the correlation lengths of magnetic fluctuations in both radial and PS runs. In the radial case, the perpendicular correlation length ($\ell_{\perp}$), which is the average of those $\hat{e}_{T}$ and $\hat{e}_{N}$, initially grows roughly proportional to $a$ as expected for uniform expansion. At later times, however, as the flow becomes magnetically dominated and quasi-two-dimensional, the co-moving $\ell_{\perp}$ begins to grow even faster than this linear $a$ scaling (Fig. 6a). This accelerated broadening of structures produces flattened, pancake-like eddies, steadily reducing $k_{\perp}$ and increasing the nonlinear time $\tau_{\rm nl}$. Visually this phase corresponds to the dominance of coherent Alfvénic vortices, growing in co-moving transverse size as the $z^{+}$ amplitude decreases. This behavior aligns with the phenomenology proposed by Meyrand2025 (Fig. 5), who argued that conservation of wave-action anastrophy—the squared magnetic vector potential in wave-action variables—drives a split transfer of energy to larger perpendicular scales. In our compressible simulations, the same qualitative behavior emerges, consistent with Alfvén vortices conserving anastrophy on intermediate timescales and expanding even faster than expansion as they decay.
@@
We also measure the critical balance parameter of the turbulent cascade using $\chi_{\rm A}\simeq(k_{\perp}z^{+})/(k_{\parallel}v_{\rm A})=(\ell_{\parallel}z^{+})/(\ell_{\perp}v_{\rm A})$. An assumption of the RDT phenomenology (discussed in section˜2.4)—that the deriving due reflection source $\mathcal{R}^{-}\tilde{z}^{+}$ balances the nonlinear sink ($\mathcal{N}^{-}\tilde{z}^{-}$) in Eq. (11)—likely requires $\chi_{\rm A}\gtrsim 1$. Only for $\chi_{\rm A}$ of order unity or larger do we expect nonlinear interactions to dominate, allowing a strong turbulent cascade to develop and damp fluctuations as assumed. In the opposite limit of $\chi_{\rm A}\gtrsim 1$ fluctuations are expected to rapidly de-correlate in $k_{\parallel}$ and drop to $\chi_{\rm A}\sim 1$ (Schekochihin2022). In Fig. 8, we thus expect the parallel scales to adjust to cause turbulence to sit at $\chi_{\rm A}\sim 1$. While there is only tentative evidence for this in the radial runs—we see a drop to $\chi_{\rm A}\approx 1.5\rightarrow 2$ initially but a larger $\chi_{\rm exp0}$ would be needed to allow a longer decay phase where this could be studied in detail—the PS runs sit near $\chi_{\rm A}\sim 1$ at late times, consistent with sustained, critically balanced turbulence. The radial runs trend toward a true 2D state with $k_{\parallel}=0$ ($\ell_{\parallel}\rightarrow 0$) at late times, but the measured $\ell_{\parallel}$ becomes box-limited (2D modes are not well captured by method) and the plotted $\chi_{\rm A}$ is therefore not necessarily a useful measure.
@@ 4.4 Fluctuation Spectra
Figure 10: Perpendicular energy spectra in the HR-$\Phi_{0}\!=\!0^{\circ}$ and HR-$\Phi_{0}\!=\!4^{\circ}$ simulations. All spectra are plotted versus dimensionless perpendicular wavenumber $k_{\perp}L_{\perp}$ ($L_{\perp}$ is the comoving perpendicular box length). Panel (a): Radial field evolution of outward and inward energy Elsässer spectra $\mathcal{E}$$^{\pm}$ at the indicated expansion times, showing the development of imbalance with scale. (b) At $a=20.5$ the magnetic ($\mathcal{E}_{M}$, purple) and kinetic ($\mathcal{E}_{K}$, green) spectra for the radial run; magnetic energy dominates across the inertial range. (c) HR-$\Phi_{0}\!=\!4^{\circ}$ Elsässer spectra at $a=5.5$, demonstrating a stronger imbalance. (d) Magnetic and Kinetic energy spectra for $\Phi_{0}\!=\!4^{\circ}$ at $a=20.1$, which display comparable inertial-range slopes.
The perpendicular power spectra of outward and inward Elsässer fluctuations are shown in Fig. 10 for the HR-$\Phi_{0}\!=\!0^{\circ}$ and HR-$\Phi_{0}\!=\!4^{\circ}$ runs. Single radial and PS snapshot at $a=5$ and $a\simeq 20$ is shown. In the inertial range, the outward/inward cascade $\tilde{\mathcal{E}}^{\pm}$ follows a slope near $k_{\perp}^{-3/2}$. The Parker spiral snapshot shows nearly identical inertial-range shapes. Observational and numerical studies of perpendicular spectra and cross-helicity (e.g., Podesta2009) likewise find similar persistent slopes across a range of solar-wind conditions. At late times, both simulations evolve toward magnetically dominated spectra, with the radial run showing stronger dominance, as expected based on the discussion above. Overall, the 1D perpendicular spectra appear largely insensitive to mean-field obliquity. This underscores that geometry primarily controls the extent and persistence of turbulence, rather than its local cascade law.
These results also highlight the limitations of 1D spectral diagnostics. The apparent similarity of the $k_{\perp}$ spectra belies the more intricate, scale-dependent anisotropy evident in Fig. 6, shaped by the spiral geometry. As observed by Verdini2018a, global spectra retain an approximate 5/3-5/3 slope even when local, field-aligned structure functions reveal strong scale-dependent anisotropy. More detailed analyses—beyond 1D average spectra—could reveal whether the Parker spiral imprints distinctive anisotropy or phase structure below the outer scale. For now, the spectra simply confirm that both runs achieve strong Alfvénic turbulence with nearly Kolmogorov-like inertial ranges (GS95).
@@ 5.1 Joint Evolution of Cross Helicity and Residual Energy
such that fully aligned fluctuations with $\sigma_{\theta}=1$ lie on the circle $\sigma^{2}_{r}+\sigma^{2}_{c}=1$. Observations show fluctuations clustered near the lower-right quadrant of the unit circle, shifting from $\sigma_{c}\!\sim\!1,\ \sigma_{r}\!\sim\!0$ close to the Sun to $\sigma_{c}\!\sim\!0,\ \sigma_{r}\!\sim\!-1$ at larger radii (Bruno2007; Bruno2013; Wicks2013). Figure 11 illustrates the trajectory of $(\sigma_{c},\sigma_{r})$ as the solar wind expands (parametrized by the expansion factor $a$), for both the pure radial field case and the PS case.
@@
In both cases, the system begins highly imbalanced ($\sigma_{c}\approx 1$, i.e. almost pure $z^{+}$) with $\sigma_{r}\approx 0$. As $a$ increases, the curve stays close to the edge of the unit circle meaning $\sigma_{r}$ becomes increasingly negative. Physically, outward $\bm{z}^{+}$ fluctuations are reflected by the inhomogeneous wind into inward $\bm{z}^{-}$ that are nearly anti-aligned ($\bm{z}^{-}\approx-\bm{z}^{+}$), so this naturally generates turbulence with $\langle\bm{z}^{-}\cdot\bm{z}^{+}\rangle<0$ (hence $\sigma_{r}<0$, Eq. 24). This eventually produces magnetically dominated Alfvénic vortices (Fig. 3) as $\chi_{\rm exp}$ decreases and $\sigma_{c}$ begins to decline toward zero (balanced turbulence), with $\sigma_{r}$ near its maximal negative values. The results are consistent with RMHD simulations of Meyrand2025 and in situ wind observations, which is transitions from imbalanced to nearly balanced turbulence with magnetic dominance beyond 1 AU (Bruno2007).

### v2

@@ 4.2 Energy Evolution and Heating
Figure 5 shows the normalized wave action energies, $\tilde{E}^{\pm}/\tilde{E}^{+}_{0}$, as functions of expansion factor $a$. Here $\tilde{E}^{+}_{0}=\tilde{E}^{+}(a=1)$ is the initial outward wave-action energy, so its decrease provides a measure of turbulent heating per (16). Plotted are HR-$\Phi_{0}\!=\!0^{\circ},4^{\circ}$; A05-$\Phi_{0}\!=\!0^{\circ},2^{\circ},5^{\circ}$, and MR-$\Phi_{0}\!=\!0^{\circ},2^{\circ},5^{\circ}$ runs spanning initial $\chi_{\rm exp,0}$ in [15, 45][$\approx$ 15,\;45]. For runs with smaller $\chi_{\rm exp,0}$ we find an initial approximate power-law decay $\tilde{E}^{+}\propto a^{-0.6}$ or $a^{-0.5}$; for larger $\chi_{\rm exp,0}$ the decay is faster with $\sim a^{-0.8}$. This discrepancy with the $a^{-1}$ limit predicted in RDT phenomenology of section˜2.4 (Eq. 15) is likely because the $a^{-1}$ decay is recovered only for sufficiently large $\chi_{\rm exp,0}$. A similar conclusion was found in Meyrand et al. (2025), who found that very large $\chi_{\rm exp}$ was needed to recover the expected $a^{-1}$ decay.
Focusing on the HR runs, the radial case displays a slow change near $a\!\sim\!15$ where $\tilde{E}^{+}$ stops decaying. We will show below that this is consistent with $\chi_{\rm exp}$ approaching unity, as discussed in the section 2. Concurrently $\tilde{E}^{-}$ grows with expansion, following a similar trend, and approaching $\tilde{E}^{+}$ as the system moves towards a quasi-balanced state. During this transition from a strongly imbalanced turbulent state to a quasi balanced sate, the nonlinear coupling weakens because $\bm{z}^{+}$ and $\bm{z}^{-}$ becomes proportional (see Fig. 3 and 4) and the turbulent heating rate drops. Eventually, as shown by Meyrand et al. (2025) we expect $\tilde{E}^{+}$ to rise as the system starts developing a very large negative residual energy,
@@
The simulations also allow a cursory examination of the influence of fluctuation amplitude on the energy evolution. Figure 5 compares (a) a high-amplitude case with $\rm A=1.0$ and (b) a lower amplitude case with $\rm A=0.5$ and similar $\chi_{\rm exp,0}$. In the lower amplitude case, the fluctuations remain nearly linearly polarized (reduced-MHD like), and follow a slower energy decay $\tilde{E}^{+}\propto a^{-0.5}$, while higher-amplitude runs, which must have larger $\delta B_{\parallel}$ to maintain spherical polarization exhibit steeper decay. One possible mechanism to explain this difference is that, the anisotropic reflection term (see Eq. 9) that seeds $\tilde{\bm{z}}^{-}$ from $\tilde{\bm{z}}^{+}$ modifies the character of the reflection once fluctuations become large. For small-amplitude, nearly transverse Alfvénic fluctuations, the reflection operator $\tilde{\mathbb{T}}=\mathrm{diag}(-1,1,1)$ leads to $\tilde{z}^{-}_{y,z}\propto-\tilde{z}^{+}_{y,z}$ with $\tilde{z}^{\pm}_{x}$ negligible. At large amplitude, however, finite $\delta B_{\parallel}$ (nonzero $\tilde{z}_{x}^{\pm}$) and the oblique mean field mix components: the 1-1 in $\tilde{\mathbb{T}}$ produces additional contributions to $\tilde{z}_{x}^{-}\propto\tilde{z}_{x}^{+}$. In other words, high amplitude fluctuations scramble the simple $\tilde{\bm{z}}^{-}\propto-\tilde{\bm{z}}^{+}$ relation, reduce the magnitude of $\langle\tilde{\bm{z}}^{+}\!\cdot\!\tilde{\bm{z}}^{-}\rangle$, and thereby allow stronger nonlinearity to persist in the PS geometry compared with the radial case. In-situ near-Sun measurements often show large normalized fluctuation amplitudes (Bale et al. 2019), the high-amplitude runs are likely more representative of observed solar wind.
@@ 4.3 Evolution of length-scales and $\chi_{exp}$
Figure 6: Evolution of correlation length $\ell$ as a function of $a$ for magnetic-field ($\bm{B_{\perp}}$) fluctuations for MR-$\Phi_{0}\!=\!0^{\circ}$ and MR-$\Phi_{0}\!=\!2^{\circ}$ runs. Each curve represents projections along the field ($\parallel$, blue), perpendicular($\perp$, brown), $\hat{e}_{T}$ direction ($\perp,T$, green), and $\hat{e}_{N}$ direction ($\perp,N$, orange). The dashed black line indicates the $a^{1}$ power-law scaling for reference. In the radial case (left panel), $\ell_{\perp}$ grows nearly linearly with $a$, consistent with expansion-driven eddy widening at early times then transitions to faster growth at late times (see Meyrand et al. 2025). In contrast, the Parker spiral geometry modifies the growth: $\ell_{\perp,T}$ saturates and remains nearly constant, while $\ell_{\perp,N}$ increases more slowly than linear.
Figure 6 shows the correlation lengths of magnetic fluctuations in both radial and PS runs. In the radial case, the perpendicular correlation length ($\ell_{\perp}$), which is the average of those $\hat{e}_{T}$ and $\hat{e}_{N}$, initially grows roughly proportional to $a$ as expected for uniform expansion. At later times, however, as the flow becomes magnetically dominated and quasi-two-dimensional, the co-moving $\ell_{\perp}$ begins to grow even faster than this linear $a$ scaling (Fig. 6a). This accelerated broadening of structures produces flattened, pancake-like eddies, steadily reducing $k_{\perp}$ and increasing the nonlinear time $\tau_{\rm nl}$. Visually this phase corresponds to the dominance of coherent Alfvénic vortices, growing in co-moving transverse size as the $z^{+}$ amplitude decreases. This behavior aligns with the phenomenology proposed by Meyrand et al. (2025) (Fig. 5), who argued that conservation of wave-action anastrophy—the squared magnetic vector potential in wave-action variables—drives a split transfer of energy to larger perpendicular scales. In our compressible simulations, the same qualitative behavior emerges, consistent with Alfvén vortices conserving anastrophy on intermediate timescales and expanding even faster than expansion as they decay.
@@
We also measure the critical balance parameter of the turbulent cascade using $\chi_{\rm A}\simeq(k_{\perp}z^{+})/(k_{\parallel}v_{\rm A})=(\ell_{\parallel}z^{+})/(\ell_{\perp}v_{\rm A})$. An assumption of the RDT phenomenology (discussed in section˜2.4)—that the deriving due reflection source $\mathcal{R}^{-}\tilde{z}^{+}$ balances the nonlinear sink ($\mathcal{N}^{-}\tilde{z}^{-}$) in Eq. (11)—likely requires $\chi_{\rm A}\gtrsim 1$. Only for $\chi_{\rm A}$ of order unity or larger do we expect nonlinear interactions to dominate, allowing a strong turbulent cascade to develop and damp fluctuations as assumed. In the opposite limit of $\chi_{\rm A}\gtrsim 1$ fluctuations are expected to rapidly de-correlate in $k_{\parallel}$ and drop to $\chi_{\rm A}\sim 1$ (Schekochihin, 2022). In Fig. 8, we thus expect the parallel scales to adjust to cause turbulence to sit at $\chi_{\rm A}\sim 1$. While there is only tentative evidence for this in the radial runs—we see a drop to $\chi_{\rm A}\approx 1.5\rightarrow 2$ initially but a larger $\chi_{\rm exp0}$ would be needed to allow a longer decay phase where this could be studied in detail—the PS runs sit near $\chi_{\rm A}\sim 1$ at late times, consistent with sustained, critically balanced turbulence. The radial runs trend toward a true 2D state with $k_{\parallel}=0$ ($\ell_{\parallel}\rightarrow\infty$) at late times, but the measured $\ell_{\parallel}$ becomes box-limited (2D modes are not well captured by method) and the plotted $\chi_{\rm A}$ is therefore not necessarily a useful measure.
@@ 4.4 Fluctuation Spectra
Figure 10: Perpendicular energy spectra in the HR-$\Phi_{0}\!=\!0^{\circ}$ and HR-$\Phi_{0}\!=\!4^{\circ}$ simulations. All spectra are plotted versus dimensionless perpendicular wavenumber $k_{\perp}L_{\perp}$ ($L_{\perp}$ is the co-moving perpendicular box length). Panel (a): Radial field evolution of outward and inward energy Elsässer spectra $\mathcal{E}$$^{\pm}$ at the indicated expansion times, showing the development of imbalance with scale. (b) At $a=20.5$ the magnetic ($\mathcal{E}_{M}$, purple) and kinetic ($\mathcal{E}_{K}$, green) spectra for the radial run; magnetic energy dominates across the inertial range. (c) HR-$\Phi_{0}\!=\!4^{\circ}$ Elsässer spectra at $a=5.5$, demonstrating a stronger imbalance. (d) Magnetic and Kinetic energy spectra for $\Phi_{0}\!=\!4^{\circ}$ at $a=20.1$, which display comparable inertial-range slopes.
The perpendicular power spectra of outward and inward Elsässer fluctuations are shown in Fig. 10 for the HR-$\Phi_{0}\!=\!0^{\circ}$ and HR-$\Phi_{0}\!=\!4^{\circ}$ runs. Single radial and PS snapshot at $a=5$ and $a\simeq 20$ is shown. In the inertial range, the outward/inward cascade $\tilde{\mathcal{E}}^{\pm}$ follows a slope near $k_{\perp}^{-3/2}$. The Parker spiral snapshot shows nearly identical inertial-range shapes. Observational and numerical studies of perpendicular spectra and cross-helicity (e.g., Podesta 2009) likewise find similar persistent slopes across a range of solar-wind conditions. At late times, both simulations evolve toward magnetically dominated spectra, with the radial run showing stronger dominance, as expected based on the discussion above. We also note that the dissipative range, where spectra steepen and eventually decay exponentially at the grid scale, appears similar in both geometries. This is expected, as the numerical dissipation mechanism in Athena++ operates identically regardless of the mean-field configuration. The geometric effects of the Parker spiral primarily influence the development and persistence of the outer scale of turbulence, rather than modifying the dissipation physics at the smallest resolved scales. Overall, the 1D perpendicular spectra appear largely insensitive to mean-field obliquity. This underscores that geometry primarily controls the extent and persistence of turbulence, rather than its local cascade law.
These results also highlight the limitations of 1D spectral diagnostics. The apparent similarity of the $k_{\perp}$ spectra belies the more intricate, scale-dependent anisotropy evident in Fig. 6, shaped by the spiral geometry. As observed by Verdini et al. (2018), global spectra retain an approximate 5/3-5/3 slope even when local, field-aligned structure functions reveal strong scale-dependent anisotropy. More detailed analyses—beyond 1D average spectra—could reveal whether the Parker spiral imprints distinctive anisotropy or phase structure below the outer scale. For now, the spectra simply confirm that both runs achieve strong Alfvénic turbulence with nearly Kolmogorov-like inertial ranges (Goldreich and Sridhar, 1995).
@@ 5.1 Joint Evolution of Cross Helicity and Residual Energy
such that fully aligned fluctuations with $\sigma_{\theta}=1$ lie on the circle $\sigma^{2}_{r}+\sigma^{2}_{c}=1$. Observations show fluctuations clustered near the lower-right quadrant of the unit circle, shifting from $\sigma_{c}\!\sim\!1,\ \sigma_{r}\!\sim\!0$ close to the Sun to $\sigma_{c}\!\sim\!0,\ \sigma_{r}\!\sim\!-1$ at larger radii (Bruno et al., 2007; Bruno and Carbone, 2013; Wicks et al., 2013). Figure 11 illustrates the trajectory of $(\sigma_{c},\sigma_{r})$ as the solar wind expands (parametrized by the expansion factor $a$), for both the pure radial field case and the PS case.
@@
In both cases, the system begins highly imbalanced ($\sigma_{c}\approx 1$, i.e. almost pure $z^{+}$) with $\sigma_{r}\approx 0$. As $a$ increases, the curve stays close to the edge of the unit circle meaning $\sigma_{r}$ becomes increasingly negative. Physically, outward $\bm{z}^{+}$ fluctuations are reflected by the inhomogeneous wind into inward $\bm{z}^{-}$ that are nearly anti-aligned ($\bm{z}^{-}\approx-\bm{z}^{+}$), so this naturally generates turbulence with $\langle\bm{z}^{-}\cdot\bm{z}^{+}\rangle<0$ (hence $\sigma_{r}<0$, Eq. 24). This eventually produces magnetically dominated Alfvénic vortices (Fig. 3) as $\chi_{\rm exp}$ decreases and $\sigma_{c}$ begins to decline toward zero (balanced turbulence), with $\sigma_{r}$ near its maximal negative values. The results are consistent with RMHD simulations of Meyrand et al. (2025) and in situ wind observations, which is transitions from imbalanced to nearly balanced turbulence with magnetic dominance beyond 1 AU (Bruno et al., 2007).